Find ratio of temperature of second and third week of june 2019
1
Expert's answer
2019-07-15T09:48:50-0400

In my region the temperatures in second and third week of June measured in Celsius scale were

24, 25, 28, 28, 24, 25, 26, 25, 24, 19, 12, 11, 21, 18.

The average of the second week is

(24 + 25 + 28 + 28 + 24 + 25 + 26)/7 = 25.71.

The average of the third week is

(25 + 24 + 19 + 12 + 11 + 21 + 18)/7 = 18.57.

The ratio is 18.57/25.71 = 0.72.